Suppose you have an element with a Radioactivity of 175 Rad. What will be the resulting
Radioactivity if it undergoes its a)1 st half-life? b)2 nd half-life? and c)5 th half-life?
1
Expert's answer
2020-10-21T13:19:22-0400

The radioactivity changes according to this equasion:

A=A0/2n;

A0 - initial radioactivity;

A - radioactivity after it undergoes n half-life decay


a) A=175/2=87.5 Rad;

b)A=175/22=175/4=43.75 Rad;

c)A=175/25=175/32=5.47 Rad;